### v1

@@ 5.2 Baseline Approach
As a baseline, we employ a greedy heuristic approach to allocate products to orders that is commonly used in real-world scheduling problems due to its simplicity and ease of implementation. This approach works as follows. First, the set of unfulfilled orders $O_{uf}$ are sorted by their due dates in ascending order. From this feasible set, we select the first $2n_{w}$ orders to form the candidate order set $O_{c}$. Then, the order with the earliest deadline in $O_{c}$ is selected and added to the current wave. Products are allocated to this order from inventory to fulfill while ensuring that the number of intermediate containers accessed is minimized. Next, within the remaining orders in $O_{c}$, the order that minimizes the number of intermediate containers accessed is chosen and added to the wave. This process is repeated, adding orders to the wave, until the wave is full or no more feasible orders remain. Once a wave is finalized, the wave is executed and next wave is constructed using the same steps with the latest inventory information. As a greedy method, this approach lacks foresight into future inventory arrivals and cannot take actions that are better in the long term.
5.3 Results and Discussion
@@
In this section, we present the results of our experiments comparing the baseline approach and the proposed approach. In order to account for the stochasticity in the product arrival distribution and order contents, we simulate each approach 30 times with different random seeds and report the average results. Further, to ensure a fair comparison, we use the same set of order and product arrival distributions for both approaches in each unique run.
@@
Given that our goal is to maximize the total number of orders fulfilled before their deadline, we use the percentage of orders fulfilled before their deadline without any preemptive processing as the evaluation metric. We also report the average delay in orders that were fulfilled after their deadline.
@@
Table 1: Comparison of order fulfillment performance with $\alpha=1$
Approach % Orders Fulfilled On Time Avg. Delay (days)
Baseline Approach 59.5% 9.6
Proposed Approach 81.7% 4.2
@@
Table 1 summarizes the results from the experiment, clearly demonstrating the superiority of the proposed approach. Specifically, the proposed algorithm achieved an 81.7% on-time fulfillment rate, significantly higher than the 59.5% achieved by the baseline method. Furthermore, the average delay in fulfilling orders was reduced to 4.2 days with the proposed approach, compared to 9.6 days under the baseline algorithm. This significant improvement in on-time fulfillment stems from the proposed approach’s ability to plan over a longer horizon and anticipate future inventory arrivals. Unlike the greedy baseline, our method can strategically delay fulfilling certain orders when it anticipates more efficient processing of a larger batch in the near future. For example, when faced with multiple approaching near-deadline orders, the proposed approach postpones less urgent orders to ensure capacity for the impending rush. The overall improvement in order fulfillment efficiency substantiates the efficacy of the proposed method in managing complex agricultural sortation systems.
In order to investigate how the performance of the proposed approach changes with relaxed definition of a complete, feasible order, we conduct additional experiments where we vary $\alpha$, the preemptive processing parameter. Recall that $\alpha$ defines the threshold for preemptively processing incomplete, partial orders. We use the percentage of missed order deadlines, i.e., the percentage of orders that were not fulfilled before deadline, as the evaluation metric to compare the performance for different values of $\alpha$.
Figure 7: Comparison of the baseline greedy approach and the proposed approach for different values of the preemptive processing parameter $\alpha$. Higher percentage of missed order deadlines indicates lower performance. Shaded region denotes the variance in the results across different runs.
Figure 7 illustrates the results of our experiments. The results clearly show that the proposed approach outperforms the baseline greedy approach for all values of $\alpha$. When $\alpha$ is close to $0.5$, the performance gap between the two approaches is narrow due to the relaxed order completion requirements that simplify the scheduling task. However, as $\alpha$ increases, the difference in performance grows, highlighting the superiority of our proposed approach in handling more stringent fulfillment conditions. It’s important to note that while lower $\alpha$ values may improve fulfillment metrics, they introduce practical challenges such as increased freight costs and logistical complexities associated with partial order shipments.
While our proposed approach significantly outperforms the baseline, it does require increased computation time and prior knowledge of product arrival distributions. However, these factors are not significant hurdles in practice. The anytime nature of our algorithm allows it to provide improving solutions as computation time increases, and the typical duration of wave fulfillment provides ample time for computation. Moreover, the assumption of distribution knowledge is reasonable in agricultural settings, where historical data and seasonal patterns often inform reliable estimates of product arrivals.
@@ 6 Conclusion and Future Work
In this paper, we introduced an adaptive hybrid tree search algorithm to address the stochastic wave scheduling problem, with a particular focus on order fulfillment in the agricultural domain. The method synergistically blends Monte Carlo tree search with domain-specific knowledge to provide a computationally tractable solution to our problem with large state and action spaces. Our approach outperforms the baseline greedy approach in terms of order fulfillment efficiency, demonstrating its effectiveness in managing complex agricultural sortation systems. The flexibility of our approach extends beyond seed order fulfillment, making it a versatile solution that can be adapted to a wide range of problems in order fulfillment and other domains that involve sequential decision-making under uncertainty.
While this paper develops the framework for fulfillment under stochastic inventory conditions, much work remains to be done in adapting to unexpected events, such as equipment malfunctions or sudden shifts in demand. Another possible area of future work is handling the interplay between the wave scheduling policy and the inventory management policy that creates products from raw material inventory. Relevant further study also includes developing hierarchical policies that plan in the space of order scheduling primitives instead of individual orders.

### v2

@@ 5.2 Baseline Approaches
The first baseline is a greedy heuristic approach (GH) to allocate products to orders that is commonly used in real-world scheduling problems due to its simplicity and ease of implementation. This approach works as follows: first, the set of unfulfilled orders $O_{uf}$ are sorted by their due dates in ascending order. From this feasible set, we select the first $2n_{w}$ orders to form the candidate order set $O_{c}$. Then, the order with the earliest deadline in $O_{c}$ is selected and added to the current wave. Products are allocated to this order from inventory to fulfill while ensuring that the number of intermediate containers accessed is minimized. Next, within the remaining orders in $O_{c}$, the order that minimizes the number of intermediate containers accessed is chosen and added to the wave. This process is repeated, adding orders to the wave, until the wave is full or no more feasible orders remain. Once a wave is finalized, the wave is executed and next wave is constructed using the same steps with the latest inventory information. As a greedy method, this approach lacks foresight into future inventory arrivals and cannot take actions that are better in the long term.
Rolling Horizon Heuristic
The Rolling Horizon Heuristic (RHH) represents a more sophisticated approach that incorporates lookahead capability into the scheduling process. At each decision point, RHH considers a fixed planning horizon that spans multiple future waves or a specified time window. Within this horizon, it attempts to optimize a sequence of waves through a multi-step process that begins with forecasting expected inventory arrivals based on historical arrival patterns. The heuristic then generates a set of candidate waves for the entire window, initially ordering them based on order deadlines.
The core of RHH’s optimization process involves selecting waves across the limited horizon to minimize a comprehensive objective function. This function accounts for multiple factors: the predicted delays for orders included in these waves, penalties for not selecting near-deadline orders that could feasibly be processed within the window, and the estimated efficiency of container access for each potential wave configuration. After determining the optimal sequence of waves for the current horizon, RHH executes only the first wave from this sequence before advancing time and repeating the entire process with a new planning horizon.
While RHH incorporates a degree of planning and anticipation of future states that makes it more sophisticated than the purely reactive GH approach, it lacks the extensive stochastic lookahead and adaptive learning of our MCTS-based proposed approach.
Vanilla MCTS Baseline
To validate the effectiveness of our domain-specific enhancements, we implement a vanilla MCTS baseline that uses standard UCT without our adaptive action space reduction mechanisms. Given the computational intractability of exploring the full action space, this baseline employs random sampling to maintain feasibility. Specifically, at each decision step, the vanilla MCTS randomly samples 1000 feasible waves, then applies the standard UCT algorithm to this randomly selected action set.
This implementation uses the same UCT formula, simulation budget, and computational resources as our proposed approach, with the key difference being the absence of intelligent action curation. Instead of our dual-criteria candidate order selection, dynamic peak reducing factor, and estimated fulfillment time filtering, the vanilla MCTS relies solely on random sampling to reduce the action space to a manageable size. This design choice allows us to isolate and quantify the impact of our domain-specific enhancements while maintaining computational tractability.
5.3 Performance Results
@@
This section presents the main experimental results comparing the performance of our proposed approach with the baseline methods. In order to account for the stochastic nature of the problem, we report statistics over 30 independent runs. Further, to ensure a fair comparison, we use the same set of order and product arrival distributions for both approaches in each unique run.
@@
Given that our goal is to maximize the total number of orders fulfilled before their deadline, we use the percentage of orders fulfilled before their deadline without any preemptive processing ($\alpha=1$) as the evaluation metric. We also report the delay statistics in orders that were fulfilled after their deadline.
@@
Table 1: Comparison of order fulfillment performance with $\alpha=1$ calculated over 30 independent runs. Delay statistics are calculated for delayed orders only.
Approach % Orders Fulfilled Avg. Delay Median Delay
On Time ($\pm$ std) (days, $\pm$ std) (days, $\pm$ std)
Greedy Heuristic (GH) 59.5% ($\pm$ 2.5) 9.6 ($\pm$ 1.2) 8.5 ($\pm$ 1.0)
Vanilla MCTS 64.3% ($\pm$ 2.3) 8.7 ($\pm$ 1.1) 7.9 ($\pm$ 0.9)
Rolling Horizon (RHH) 68.2% ($\pm$ 2.2) 7.5 ($\pm$ 1.0) 6.8 ($\pm$ 0.8)
Proposed Approach 81.7% ($\pm$ 1.8) 4.2 ($\pm$ 0.8) 3.5 ($\pm$ 0.6)
Table 1 summarizes the results from the experiment, clearly demonstrating the superiority of our proposed approach across all metrics. Our method achieved an 81.7% on-time fulfillment rate, substantially higher than the 59.5% achieved by GH, the 64.3% achieved by vanilla MCTS, and the 68.2% by RHH. The performance advantage extends to delay metrics as well, with our approach reducing the average delay for late orders to just 4.2 days, compared to 9.6 days with GH, 8.7 days with vanilla MCTS, and 7.5 days with RHH.
The vanilla MCTS results are particularly revealing. Despite using the same computational budget and UCT algorithm as our proposed approach, the vanilla MCTS achieves only marginal improvements over the simple greedy heuristic. This modest 4.8 percentage point improvement (from 59.5% to 64.3%) demonstrates that MCTS alone, without domain-specific guidance, struggles to effectively navigate the vast action space. The random sampling of waves, while computationally necessary, fails to consistently identify high-quality scheduling decisions, resulting in performance that falls well short of both RHH and our proposed method.
@@
This significant performance improvement stems from the proposed approach’s ability to effectively plan over longer horizons and anticipate future inventory arrivals. Unlike the baselines, our approach can strategically delay processing certain orders when it anticipates more efficient batch processing opportunities in the near future or when it detects potential bottlenecks ahead. For example, when faced with multiple approaching deadline clusters, our method can sacrifice processing less urgent orders to ensure capacity for an impending rush of near-deadline orders.
The performance gap between RHH and GH confirms the value of even limited lookahead planning. More importantly, the substantial 17.4 percentage point improvement of our approach over vanilla MCTS conclusively demonstrates that intelligent action space curation is essential for making MCTS effective in this domain.
5.4 Sensitivity Analysis
To examine the robustness of our approach across different operational conditions, we conducted sensitivity analyses varying key parameters: the preemptive processing parameter ($\alpha$) and three warehouse constraints (container capacity, induction stations, and number of chutes/maximum wave size).
Figure 7: Sensitivity analysis of key parameters: (a) preemptive processing parameter ($\alpha$), (b) container capacity, (c) number of induction stations, and (d) number of chutes (maximum wave size). For the preemptive processing parameter, lower missed order rates indicate better performance. For the warehouse constraints, higher on-time fulfillment percentages indicate better performance. All metrics are visualized using line plots with distinctive markers for each approach (circle for GH, square for RHH, triangle for proposed method). Consistent coloring is used across all panels: red for Greedy Heuristic, green for Rolling Horizon Heuristic, and blue for our proposed approach.
Effect of Preemptive Processing Parameter
Figure 7a shows the percentage of missed order deadlines across different values of $\alpha$ for all three approaches. Lower values of $\alpha$ allow processing orders with partial inventory when the deadline is approaching, while higher values require more complete inventory availability before processing.
The results demonstrate that our proposed approach consistently outperforms both baseline methods across all values of $\alpha$, with the performance gap widening as $\alpha$ increases. When $\alpha$ approaches 0.5, the performance gap narrows because the relaxed order completion requirements simplify the scheduling task. However, as $\alpha$ increases toward 1.0 (requiring full order completion), the scheduling problem becomes more challenging, and the sophisticated planning of our approach shows greater benefits. Statistical tests confirm that the performance differences are significant at all tested $\alpha$ values.
Effect of Warehouse Constraints
Figures 7b, 7c, and 7d present the impact of varying warehouse physical constraints on order fulfillment performance across extended parameter ranges. These constraints directly affect the operational capacity and efficiency of the fulfillment system.
Container Capacity
Figure 7b shows that increasing container capacity from 150 to 350 items consistently improves performance across all approaches, as larger containers reduce the total number of containers needed and enhance picking efficiency. Our proposed approach shows consistent superiority, maintaining approximately a 20 percentage point advantage in on-time fulfillment over GH and a 13 point advantage over RHH across all capacity settings. The performance gains show a diminishing returns pattern at higher capacities, suggesting that beyond a certain threshold, other system constraints become limiting factors.
Induction Stations
Figure 7c illustrates how the number of induction stations affects how quickly items can be introduced into the sortation system. While increasing the number of stations improves performance for all methods, our approach demonstrates robust performance even under constrained conditions, achieving 72.5% on-time fulfillment with only 10 stations compared to GH’s 46.5% and RHH’s 55.4%.
Number of Chutes
Figure 7d shows how the number of chutes, which directly affects the maximum wave size ($n_{w}$), impacts performance across a range from 200 to 600 chutes. Increasing this parameter yields moderate performance improvements across all approaches, with our method maintaining its significant advantage throughout the range. This indicates that while larger waves can improve efficiency to some extent, intelligent wave formation and scheduling are more critical factors in performance.
Overall, these sensitivity analyses demonstrate that our approach’s advantage is robust across various operational configurations and parameter ranges. Even under the most constrained conditions tested, the long-horizon planning and adaptive nature of our method maintain a substantial performance edge over both baseline approaches. The consistent pattern across all parameter variations highlights the fundamental superiority of our approach’s planning capabilities rather than its reliance on specific operational conditions.
5.5 Scheduling Performance by Order Urgency
In order to provide more granular insight into how scheduling decisions impact different types of orders, we analyzed fulfillment statistics for orders categorized by their initial deadline proximity when first considered by the scheduler. This analysis helps understand how each approach handles trade-offs between urgent and less time-sensitive orders.
Table 2: Order fulfillment performance by initial deadline proximity (30 runs, $\alpha=1$). Performance shown as percentage of orders fulfilled on time and average delay in days.
Deadline Greedy Heuristic Vanilla MCTS Rolling Horizon Proposed
Urgent 45.2% / 5.1d 52.1% / 4.3d 58.4% / 3.5d 75.8% / 2.2d
Medium 60.5% / 9.2d 65.3% / 8.5d 69.1% / 7.0d 82.3% / 4.1d
Low 72.8% / 12.5d 75.5% / 11.2d 78.1% / 9.8d 87.0% / 6.3d
Table 2 presents performance metrics across three deadline categories: Urgent (deadline << 7 days), Medium Priority (deadline 7-21 days), and Low Priority (deadline >> 21 days). All approaches demonstrate better on-time performance for less urgent orders due to increased scheduling flexibility, with our proposed approach achieving superior results across all categories.
The most significant performance difference appears in urgent order handling, where our approach achieves 75.8% on-time fulfillment compared to 45.2% for GH and 58.4% for RHH. This demonstrates our method’s effectiveness in prioritizing time-sensitive orders. The average delays are also consistently lower with our approach - for urgent orders, we reduce delays to 2.2 days versus 5.1 and 3.5 days for GH and RHH respectively.
Our approach shows more balanced performance across order categories, with only a 11.2 percentage point difference between low-priority and urgent orders. While GH’s simple deadline-based approach struggles with urgent orders when resources are misaligned, and RHH shows limited improvement with its short planning horizon, our proposed method excels by balancing priorities between urgent and less critical orders, achieving superior performance across all categories.
5.6 Computational Considerations
Real-time warehouse applications necessitate careful consideration of computational demands for the proposed approach. The MCTS algorithm’s complexity per decision step is approximately $N\times D_{avg}\times C_{s}$, where $N$ is the simulation count, $D_{avg}$ is their average depth, and $C_{s}$ is the simulation step cost. Our domain-specific heuristics for candidate order selection (sorting, $O(|O_{uf}|\log|O_{uf}|)$) and action generation (sampling and filtering) are designed for efficiency, managing $C_{s}$ and reducing the MCTS branching factor.
Table 3: Average computation time per decision step with standard deviations.
Algorithm Avg. Time per Decision (sec)
Greedy Heuristic (GH) 0.42 ± 0.08
Vanilla MCTS 10.85 ± 2.11
Rolling Horizon Heuristic (RHH) 3.15 ± 0.85
Proposed Approach (MCTS) 11.37 ± 2.24
Table 3 presents empirical computation times for a simple warehouse scenario. The Greedy Heuristic (GH) exhibits very fast and consistent performance (0.42 ± 0.08 sec), the Rolling Horizon Heuristic (RHH) shows moderate variability (3.15 ± 0.85 sec), while our approach requires longer but still practical computation time (11.37 ± 2.24 sec), configurable by adjusting the MCTS simulation budget.
Despite this higher computational cost of our approach, practical factors support its viability. Wave scheduling decisions typically occur at larger time scales, and MCTS’s anytime property permits flexible interruption. Crucially, the substantial performance gains (e.g., $20\%$ improvement in on-time fulfillment) often justify this cost, as late deliveries are expensive. Empirical tests confirm tractability in high-throughput settings with adequate inter-wave planning time.
@@ 6 Conclusion and Future Work
In this paper, we introduced an adaptive hybrid tree search algorithm to address the stochastic wave scheduling problem, with a particular focus on order fulfillment in the agricultural domain. The method blends Monte Carlo tree search with domain-specific knowledge to provide a computationally tractable solution to our problem with large state and action spaces. Our approach outperforms the baseline greedy approach in terms of order fulfillment efficiency, demonstrating its effectiveness in managing complex agricultural sortation systems. The flexibility of our approach extends beyond seed order fulfillment, making it a versatile solution that can be adapted to a wide range of problems in order fulfillment and other domains that involve sequential decision-making under uncertainty.
While the presented results are promising, it is important to acknowledge certain limitations of the current study. Our validation relies on simulation, and while our simulator closely mirrors real operations, testing in an actual agricultural fulfillment center remains necessary to verify practical effectiveness. Additionally, the product arrival model’s accuracy depends heavily on historical data quality, making our approach vulnerable to unforeseen shifts in arrival patterns.
Building on this work, several avenues for future research warrant exploration. One key direction is enhancing real-time adaptability; this includes developing an online learning mechanism for the peak reducing factor $\rho$ using a Bayesian update scheme, enabling within-season adaptation to improve performance during atypical arrival patterns. Additionally, hierarchical action abstraction through order clustering could significantly reduce computational complexity and enable longer planning horizons. Finally, improving the robustness of the product arrival model to non-stationarity is important; this could involve researching techniques to make the TVMC more quickly adaptable to shifts in seasonal patterns not present in historical data, or integrating online learning mechanisms.